40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 845     Entry time: Mon Aug 18 09:19:55 2008
Author: steve 
Type: Summary 
Category: VAC 
Subject: 11 days at atm 
It took 11 days to fix earth quake triggered sus problems of ITMX, SRM and PRM
Only ITMX north and BSC north vac doors were removed.

The PRM sus had to be removed form the vac envelope for "hip replacement"-new wire stand off was
epoxied in place.
Note: the PRM has no guide rod on the other side

ITMX, SRM and BS osems were optimized in place.
No crosscoupling optimization was performed.
Beam block was removed from ITMXC, it was too close to the main beam.

POX pick off mirror and mount will be replaced next vent.

Vac viewports were inspected from the inside.
Attachment 1: ventprm.jpg  85 kB  | Hide | Hide all
ventprm.jpg
ELOG V3.1.3-